# Perverse microsheaves

### v2

@@ 4.4. Maslov data
By $C$-Maslov data for $V$, we mean a null-homotopy of the map $V\to B^{2}Pic(C)$.
A choice $\eta$ of $C$-Maslov data gives a trivialization of the $BPic(C)$ bundle to which Theorem 4.14 asserts that $\mu sh_{V^{LGr}}$ descends. We write $\mu sh_{V,\eta}$ for the sheaf of categories on $V$ obtained by the pullback along the zero section of this trivial bundle (of groups).
@@
For a subset $X\subset V$ we write $\mu sh_{X,\eta}\subset\mu sh_{V,\eta}$ for the sheaf of full subcategories on objects supported in $X$. For any smooth Legendrian $L\subset V$, Theorem 4.14 plus the Weinstein neighborhood theorem imply $\mu sh_{L,\eta}\subset\mu sh_{V,\eta}$ depends only on $\eta|_{L}$ (and not on $V$).
@@
There is a canonical choice of Maslov data on (or in a neighborhood of) a real Legendrian $L$. Indeed, canonically $\xi|_{L}\cong T^{*}L$, and taking the fiber polarization $\phi_{L}$ gives a canonical null-homotopy of $L\to V\to BLGr$. By composition we obtain a canonical null-homotopy $\phi_{L}$ of $L\to B^{2}Pic(C)$.
Given Maslov data $\eta$ for $V$, we obtain another such null-homotopy $\eta|_{L}$, we say that a homotopy $\phi_{L}\sim\eta|_{L}$ is secondary Maslov data for $L$. Evidently, secondary Maslov data induces an isomorphism $\mu sh_{L,\eta|_{L}}\cong\mu sh_{L,\phi_{L}}$.
It is possible to further fix an isomorphism $\mu sh_{L,\phi_{L}}\cong Loc_{L}$, the right hand side being the sheaf of categories of local systems on $L$. To do so, consider the embedding $L=L\times 0\subset L\times\mathbb{R}$. Let $L^{+}\subset S^{*}(L\times\mathbb{R})$ be the positive conormal to $L$.
Observe that the fiber polarization of $S^{*}(L\times\mathbb{R})$ restricts to $\phi_{L^{+}}$. Taking the pushforward along the inclusion and then microlocalizing along the positive conormal gives a map $Loc_{L}\to\mu sh_{L}\cong\mu sh_{L,\phi_{L}}$, which we can check locally is an equivalence (e.g. using [15, Proposition 6.6.1]).
Remark 4.17.
The category of co-orientable contact manifolds is equivalent to the category of symplectic manifolds with a free transitive $\mathbb{R}_{>0}$ action scaling the symplectic form: given a contact manifold $V$, we may form the corresponding symplectic manifold $\mathbb{R}_{>0}V$ by taking the locus in $T^{*}V$ of all graphs of contact 1-forms respecting the co-orientation. An embedding $V\hookrightarrow S^{*}M$ always lifts to a $\mathbb{R}_{>0}$-equivariant embedding $\mathbb{R}_{>0}V\to T^{\circ}M=\mathbb{R}_{>0}S^{*}M$, uniquely up to shift by a function $f:V\to\mathbb{R}_{>0}$, and a thickening of $V$ canonically lifts to a thickening of $\mathbb{R}_{>0}V$, etc. Stable polarizations and Maslov data being topological, these notions are equivalent on $V$ and $\mathbb{R}_{>0}V$.
Thus by embedding, thickening, descending, etc., we find that from Maslov data $\eta$ on $V$, we obtain an $\mathbb{R}_{>0}$-equivariant sheaf of categories $\mu sh_{\mathbb{R}_{>0}V,\eta}$ on $\mathbb{R}_{>0}V$. It is of course canonically isomorphic to the pullback of $\mu sh_{V,\eta}$ along the projection $\mathbb{R}_{>0}V\to V$.
In the discussion thus far, we have been agnostic as far as the choice of the category $C$, and we have also not needed to compute the map $\mu_{C}:U/O\to BPic(C)$. We now turn to this question. Note first that given a map of symmetric monoidal stable categories $C\to D$, it is clear that $\mu_{D}$ is the composition of $\mu_{C}$ with the natural map $Pic(C)\to Pic(D)$.
In particular, when $R$ is an ordinary commutative ring ($R=\pi_{0}(R)$), the map $\mu_{R-mod}$ factors through $\mu_{\mathbb{Z}-mod}$. We have $Pic(\mathbb{Z}-mod)=\mathbb{Z}\oplus B(\mathbb{Z}/2\mathbb{Z})$, where $\mathbb{Z}/2\mathbb{Z}$ is really the group of invertible integers, {±$\{\pm 1\}$. The map $\mu_{\mathbb{Z}-mod}$ has been calculated:
Theorem 4.18.
[8] The map $\mu_{\mathbb{Z}-mod}$ of Theorem 4.14 is $LGr\xrightarrow{\det^{2}\oplus w_{2}}S^{1}\oplus B^{2}(\mathbb{Z}/2\mathbb{Z})$. That is, $\mathbb{Z}-mod$ Maslov data is precisely grading and orientation data.
It follows that grading and orientation data provide $R-mod$ Maslov data for any commutative ring $R$, although of course there may be $R-mod$ Maslov data which does not arise in this way.
Remark 4.19.
Let $C=\mathbb{Z}-mod$. Let $M$ be a manifold. Let $m_{can}$ be the Maslov datum whose grading comes from the fiber polarization, and whose orientation data is the canonical orientation data of Def. 3.8. By Lemma 3.9, Corollary 4.13, and Theorem 4.18, a spin structure on $M$ gives an equivalence of sheaves of categories on $T^{\circ}M$: $\mu sh_{S^{*}M}\cong\mu sh_{S^{*}M,m_{can}}$ where the right hand side is defined by Equation (9) and the LHS by Theorem 4.14.
If $M=N\times\mathbb{R}$ is not spin, then in fact $\mu sh_{S^{*}M}\not\cong\mu sh_{S^{*}M,m_{can}}$. Indeed, in this case the conormal Legendrian to $N\times 0$ can be seen to support a microsheaf in $\mu sh_{S^{*}M}$ but not in $\mu sh_{S^{*}M,m_{can}}$.
This corresponds to the observation of Kragh that $Fuk(T^{*}M)\neq Loc(M)$ when $M$ is not spin [17, Remark 1.3].
Remark 4.20.
We return to the case of general $C$. Consider first the universal case $C=\mathbb{S}-mod$ of spectra. The map $\mu_{\mathbb{S}-mod}$ was calculated in [9]; the result is that it is a delooping of the $J$-homomorphism [9]. The map $\mathbb{S}\to\pi_{0}(\mathbb{S})=\mathbb{Z}$ induces $Pic(\mathbb{S}-mod)\to\pi_{\leq 1}Pic(\mathbb{S}-mod)=Pic(\mathbb{Z}-mod)$. In particular, the inclusion of the shift $\mathbb{Z}\to Pic(\mathbb{S}-mod)$ splits and we can write $Pic(\mathbb{S}-mod)=\mathbb{Z}\oplus Pic(\mathbb{S})_{0}$, where $Pic(\mathbb{S})_{0}$ is the identity connected component. (It is known that $Pic(\mathbb{S})_{0}=B\mathbb{S}^{\times}$ for some $\mathbb{S}^{\times}$ with a natural interpretation as “the units of $\mathbb{S}$”.) We write $w_{\mathbb{S}-mod}$ for the composition $LGr\to BPic(\mathbb{S}-mod)\to B^{2}\mathbb{S}^{\times}$, and define $\mathbb{S}-mod$ orientation data to be a null-homotopy of the appropriate composition with $Bw_{\mathbb{S}-mod}$. Tautologically, then, $\mathbb{S}$-mod Maslov data is the same as a grading and $\mathbb{S}-mod$ orientation data.
For more general $C$, let us write $Pic(C)_{0}$ for the identity connected component. Then the canonical map $\mathbb{S}-mod\to C$ induces a map $Pic(\mathbb{S})_{0}\to Pic(C)_{0}$, and we write $w_{C}$ for the composition of $w_{\mathbb{S}-mod}$ with the delooping of this map. We define $C$-orientation data to be a null-homotopy of the appropriate composition with $Bw_{C}$. Evidently a grading and $C$-orientation datum yield a $C$-Maslov datum, though not all $C$-Maslov data need arise in this way.
Still more generally, suppose given any partition $\Pi$ of the set of isomorphism classes of objects of $C$. We write $Pic(C)_{\Pi}\subset Pic(C)$ for the full subcategory on invertible elements which preserve the partition $\Pi$. By $(C,\Pi)$-orientation data, we mean a null-homotopy of the composite map to $BPic(C)_{\Pi}$. ($C$-orientation data being the case of the total partition). In particular, for $D\subset C$ a subcategory, we consider the partition into objects in versus not in $D$, and term the corresponding orientation data as $(C,D)$-orientation data.
5. Microsheaves on complex contact manifolds

### v3

@@ A $(\mathcal{C},\{\mathcal{D}_{i}\})$-grading is a null-homotopy of the map
We can give “constrained” analogues of the constructions in Section 4.6. Namely, suppose now that $L\subset V$ is a Legendrian. If $\{\mathcal{D}_{i}\}$ is a collection of subcategories, as in Definition 4.34, a secondary orientation on $L$ constrained by $\{\mathcal{D}_{i}\}$ is a homotopy between the polarization $(\mathcal{C},\{\mathcal{D}_{i}\})$-grading and the $(\mathcal{C},\{\mathcal{D}_{i}\})$-grading induced from $\eta$.
A secondary Maslov datum for $L$ lifting a given secondary polarization constrained by $\{\mathcal{D}_{i}\}$ shall be called a secondary orientation datum for $L$ constrained by $\{\mathcal{D}_{i}\}$.
Corollary 4.36 (cf. Corollary 4.30).
Let $\eta$ be a Maslov datum and $L$ be an Legendrian endowed with a secondary grading constrained by $\{\mathcal{D}_{i}\}$. Then the ambiguity of the equivalence $\mathcal{C}\cong(\mu sh_{L,\eta})_{p}$ from Corollary 4.28 can be reduced to $\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$.
In particular, per Corollary 4.36, the statement that “the microstalk of $\mathcal{F}\in(\mu sh_{L,\eta})|_{p}$ is contained in $\mathcal{D}_{i}$” is well-defined.
Example 4.37.
Suppose that $(\mathcal{C},\mathcal{D})$ are as in Example 4.35. Then the microstalk of $\mathcal{F}\in(\mu sh_{L,\eta})|_{p}$ is an object in $\operatorname{Coh}(\mathbb{P}^{1})$ which is well-defined up to tensoring with $\mathcal{O}(n)$. Hence, it is meaningful to ask whether the microstalk of $\mathcal{F}$ belongs to the heart $\operatorname{Coh}(\mathbb{P}^{1})^{\heartsuit}$ (i.e. whether it is represented by an honest coherent sheaf).
5. Microsheaves in the complex setting
5.1. Microsheaves on complex cotangent bundles
We now review the results of Waschkies [42]. Denote by $\pi:T^{\circ}M\rightarrow\mathbb{P}^{*}M$ the projection. The perverse microsheaves on $\mathbb{P}^{*}M$ [42, Definition 6.1.2] is defined as a subsheaf of the following sheaf.
@@ 5.1. Microsheaves on complex cotangent bundles
We define the presheaf $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}^{pre}$ on $\mathbb{P}^{*}M$ by
$\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}^{pre}(\Omega)\coloneqq\mu sh^{pre}(\pi^{-1}(\Omega))$
for an open set $\Omega\subseteq\mathbb{P}^{*}M$, and denote its sheafification by $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}$. If $\Lambda\subseteq\Omega$ is a closed complex Legendrian, we can similarly define a presheaf on $\Omega$, using the notation in (13), by
$\operatorname{\mathbb{P}\mu sh}_{\Lambda}^{pre}(\Omega^{\prime})\coloneqq\mu sh_{\pi^{-1}(\Lambda)}^{pre}(\pi^{-1}(\Omega)),$
and denote its sheafification by $\operatorname{\mathbb{P}\mu sh}_{\Lambda}$. Per Lemma 4.8, the canonical map $\operatorname{\mathbb{P}\mu sh}_{\Lambda}\hookrightarrow\operatorname{\mathbb{P}\mu sh}$ is fully faithful and its image consists of those objects supported in $\Lambda$. Lastly, we define $\operatorname{\mathbb{P}\mu sh}_{\mathbb{C}-c}$ to be the subsheaf consisting of objects supported on complex Legendrians, which can also be obtained as the shaefification of $\operatorname{\mathbb{P}\mu sh}_{\mathbb{C}-c}^{pre}(\Omega)=\mu sh_{\mathbb{C}-c}^{pre}(\pi^{-1}(\Omega))$
The following Theorem 5.3 is one of the main theorems in [42], which gives a simple description of $\operatorname{\mathbb{P}\mu sh}_{\Lambda}$ at stalks when $\Lambda$ is in generic position:
Definition 5.2.
Suppose that $M$ is complex analytic and $\Lambda\subset P^{*}M$ is a (singular) complex Legendrian. Let $r:P^{*}M\to M$ be the projection. We say that $\Lambda$ is in generic position if that $m\in M$ has the property that $\Lambda\cap r^{-1}(m)=\{p_{1},\dots,p_{k}\}$ is a finite set.
Theorem 5.3 ([42, Thm. 5.1.5]).
In the situation of Definition 5.2, let $p\in r^{-1}(m)\cap\Lambda$. There is a fully faithful functor
$(\operatorname{\mathbb{P}\mu sh}_{\Lambda})_{p}=\mu sh^{pre}_{\pi^{-1}(\Lambda)}(\mathbb{C}^{*}\cdot p)\to(sh_{\mathbb{C}-c}/loc)_{m}=(sh_{\mathbb{C}-c})_{m}/loc_{m}.$
Its essential image is the full subcategory of $(sh_{\mathbb{C}-c}/loc)_{m}$ on objects whose microsupport is contained in $\Lambda\cap Op(\mathbb{C}^{*}\cdot p)$ for some small open neighborhood $Op(\mathbb{C}^{*}\cdot p)$.
Proof.
The fully faithful functor is furnished by Waschkies [42, Thm. 5.1.5] using the microlocal cutoff. Let us temporarily denote by $\mathcal{A}$ the full subcategory of $(sh_{\mathbb{C}-c}/loc)_{m}$ on objects whose microsupport is entirely contained in $\pi^{-1}(\Lambda)\cap Op(\mathbb{C}^{*}\cdot p)$. That Waschkies’ functor lands inside $\mathcal{A}$ is a consequence of [42, Thm. 5.1.5(3)]. We should prove that every $F\in\mathcal{A}$ lies in the image of Wacshkies’ functor. We may assume $M$ is a ball, so that we have coordinates $T^{\circ}M=M\times T^{\circ}_{m}M$. By assumption, there is some neighborhood $U\subset M$ of $m$ such that $SS(F)\cap\dot{\pi}^{-1}(U)\subset U\times\gamma$, where $\gamma$ is a neighborhood of $\mathbb{C}^{*}\cdot p$. Then by construction, Waschkies’ map sends (the object in $\mu sh^{pre}_{\Lambda}(\mathbb{C}^{*}\cdot p)$ represented by) $F$ to $\Phi_{U,\gamma}(\mathcal{F})$, where $\Phi_{U,\gamma}(-)$ is the microlocal cutoff. There is always a map $\alpha:\Phi_{U}(F)\to F$, and we need to show that the cone is a local system on $U$. Up to shrinking $M$, assume $U=M$. Then $cone(\alpha)$ has no microsupport in $\dot{\pi}^{-1}(M)\setminus(\gamma\times V)$ (because neither $F$ nor $\Phi_{U,\gamma}F$ have microsupport there). But $cone(\alpha)$ also has no microsupport in $U$ because $\Phi_{U,\gamma}$ induces an isomorphism in $\mu sh^{pre}(\gamma\times M)$ [42, Def. 2.3.1(iii)]. Hence the microsupport of $cone(\alpha)$ is contained in the zero section, as claimed. ∎
Remark 5.4.
The analogue of Theorem 5.3 with complex analyticity hypotheses removed and with $\mathbb{C}^{*}$ replaced by $\mathbb{R}^{+}$ is also true (indeed easier: the cutoffs already in [25] are good enough and one does not need [8]); see [35, Lem 6.7, Prop. 6.9].
Now, we compare the two notions of microsheaves $\mu sh_{T^{*}M}$ and $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}$. The canonical map
$\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}^{pre}(\Omega)\coloneqq\mu sh^{pre}(\pi^{-1}(\Omega))\rightarrow\mu sh(\pi^{-1}(\Omega))$
defines a map $\operatorname{\mathbb{P}\mu sh}_{P^{*}M}\rightarrow\pi_{*}(\mu sh|_{T^{\circ}M})$.
Lemma 5.5.
The map $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}\hookrightarrow\pi_{*}\left(\mu sh|_{T^{\circ}M}\right)$ is fully-faithful.
Proof.
It’s enough to check on stalks. Take $p\in\mathbb{P}^{*}M$, the map $(\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M})_{p}\rightarrow\left[\pi_{*}\left(\mu sh|_{T^{\circ}M}\right)\right]_{p}$ is given by
$\mu sh^{pre}(\pi^{-1}(p))\rightarrow\mu sh(\pi^{-1}(p)).$
As mentioned in (10), the Hom sheaf on the right-hand side $\mathcal{H}om_{\mu sh}$ is computed by $\mu hom$. In particular, $\operatorname{Hom}_{\mu sh(\pi^{-1}(p))}$ is computed by $\Gamma(\pi^{-1}(p);\mu hom(-,-)).$ But [42, Proposition 2.4.4] shows that it is also the case for the category on the left-hand side so we are done. ∎
Since $\mu sh^{pre}(\pi^{-1}(p))=sh(M)/sh_{T^{*}M\setminus\pi^{-1}(p)}(M)$ is a quotient, objects in $\mu sh^{pre}(\pi^{-1}(p))$ are presented by sheaves. Thus, the proof of the above lemma shows characterizes the image as the following:
Corollary 5.6.
The subsheaf $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}$ in $\pi_{*}(\mu sh_{T^{\circ}M})$ is equivalent to the subsheaf stalkwisely on $P^{*}M$ presented by a sheaf
$\{\mathcal{F}\in\pi_{*}(\mu sh|_{T^{\circ}M})|\mathcal{F}|_{p}\in\operatorname{im}(sh(M)\rightarrow\mu sh(\pi^{-1}(p)),\forall p\in\mathbb{P}^{*}M\}.$
Similar to the real situation Theorem 4.10, there is a complex version of the contact transform.
Theorem 5.7 ([25, (11.4.8)]).
Let $\mathcal{U}\subseteq\mathbb{P}^{*}M$, and $\mathcal{V}\subseteq\mathbb{P}^{*}N$ be open sets, and $\chi:\mathcal{U}\xrightarrow{\mathcal{V}}$ be a complex contactomorphism. Then, for any given $p\in\mathcal{U}$, shrink $\mathcal{U}$ if needed, one can assume that there exists a sheaf $K\in sh(M\times N)$ such that the functor $\Phi_{K}:sh(M)\rightarrow sh(N)$ given by convolving with $K$ induces an equivalence, often referred as contact transformation,
@@
where $\widetilde{\mathcal{U}}$, $\widetilde{\mathcal{V}}$ and $\tilde{\chi}$ are the corresponding symplectic lifts. Consequently, it induces an equivalence $\mu sh_{T^{\circ}M}|_{\mathcal{U}}\xrightarrow{\sim}\tilde{\chi}^{*}\mu sh_{T^{\circ}N}|_{\mathcal{V}}$ which commutes with the canonical map $\mu sh^{pre}\rightarrow\mu sh$.
Corollary 5.8.
With the notation as above, $\Phi_{K}$ induces an equivalence ${\pi_{M}}_{*}\mu sh_{\widetilde{\mathcal{U}}}\xrightarrow{\sim}\chi^{*}{\pi_{N}}_{*}\mu sh_{\widetilde{\mathcal{V}}}$ which restricts to $\operatorname{\mathbb{P}\mu sh}_{\mathcal{U}}\xrightarrow{\sim}\chi^{*}\operatorname{\mathbb{P}\mu sh}_{\mathcal{V}}$.
Proof.
The first equivalence is tautological. The second equivalence follows from the characterization of $\operatorname{\mathbb{P}\mu sh}$ in Corollary 5.6 as locally presentable by sheaves and the fact that contact transformation commutes with the canonical map $\mu sh^{pre}\rightarrow\mu sh$ as mentioned in Theorem 5.7. ∎
In general, we do not know if the inclusion $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M}\hookrightarrow\pi_{*}(\mu sh_{T^{\circ}M})$ is an equivalence. However, it is the case when we restrict to complex constructible objects.
Proposition 5.9.
The inclusion $\operatorname{\mathbb{P}\mu sh}_{\mathbb{P}^{*}M,\mathbb{C}-c}\xhookrightarrow{\sim}\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})$ is an equivalence.
Proof.
Let $p\in\mathbb{P}^{*}M$ and consider $\mathcal{F}\in[\pi_{*}(\mu sh_{T^{\circ}M,\mathbb{C}-c})]_{p}=\mu sh_{T^{\circ}M,\mathbb{C}-c}(\pi^{-1}(p))$. Since objects in $\mu sh_{T^{\circ}M,\mathbb{C}-c}(\pi^{-1}(p))$ are germs of microsheaves near $\pi^{-1}(p)$, we can pick some $\Omega\subseteq\mathbb{P}^{*}M$ containing $p$ and realize $\mathcal{F}$ as a microsheaf on $\pi^{-1}(\Omega)$ with $ss(F)\subseteq\pi^{-1}(\Omega)$ being a complex Lagrangian. Denote by $\Lambda\subseteq\Omega$ the corresponding complex Legendrian.
By the previous Corollary 5.8, being in $\operatorname{\mathbb{P}\mu sh}$ is invariant under contact transform, so we can apply the Kashiwara–Kawai general position theorem (see [23, Sec. 1.6]) and assume that the composition $\Lambda\subseteq\mathbb{P}^{*}M\xrightarrow{r}M$ is finite to one near $p$. Shrinking $\Omega$ and $M$ if needed, we may assume that $\Lambda$ admits the standard form [26, (2.5)]: There exists local coordinates $(z,\xi)$ such that $p=dz_{n}$ and $\Lambda$ is of the form $\mathbb{P}^{*}_{S}M$, where $S=\{f=0\}$ is the zero locus of some holomorphic function $f=z_{n}^{k}+g(z)$ for some $k\in\mathbb{N}$ and for some $g(z)\in(z_{1},\cdots,z_{n})^{k+1}$.
But this implies that $\mathcal{F}\in\mu sh_{\mathbb{P}_{S}^{*}M}(\Omega)$ so we can apply Theorem 4.11 and conclude that there exists an $F\in sh(M)$ which projects to $\mathcal{F}$, which in particular implies that $\mathcal{F}\in\operatorname{\mathbb{P}\mu sh}_{P^{*}M,\mathbb{C}-c}$. ∎
5.2. Canonical microsheaves and microstalks
We return to considering an exact symplectic or contact manifold $W$ with corresponding structure morphism $W\rightarrow BU$. Suppose given moreover a lift to stable quaternionic structure $W\rightarrow BSp$, e.g. arising from an underlying complex symplectic or contact structure as in Section 2. Per Definition 3.8, $W$ carries a canonical orientation/grading datum. Per Theorem 4.27, this determines a canonical $R$-Maslov datum, for $R$ a commutative ring $R$. Definition 4.25 thus furnishes a sheaf $\mu sh_{W}$ of $R$-linear categories on $W$.
More generally, fix a coefficient category $\mathcal{C}$ and an anchored collection of subcategories $\{\mathcal{D}_{i}\}$. Endow $W$ with the $(\mathcal{C},\{\mathcal{D}_{i}\})$ grading induced by the canonical $\mathbb{S}$-grading (see Definition 3.8 and (17)), and sssume given a $(\mathcal{C},\{\mathcal{D}_{i}\})$ orientation $o$. Then Definition 4.25 furnishes a sheaf of $\mathcal{C}$-linear categories on $W$ which we denote by $\mu sh_{W,o}$.
Proof of Theorem 1.1.
Let us first discuss the case $R=\mathbb{Z}$. We take the canonical $\mathbb{Z}$-Maslov datum. Assume $L$ is a complex Lagrangian or Legendrian. Then, the polarization $\phi_{L}:L\rightarrow BO$ admits a lift to a complex polarization $L\rightarrow BU\rightarrow BO$. To compare it to the fiber polarization $\phi_{L}$ near L, we consider the secondary Maslov data. By Theorem 4.27 secondary Maslov data are in one-to-one correspondence with secondary orientation/grading. The theorem thus reduces to Lemma 3.13.
For general $R$, note that Lemma 3.13 is itself deduced from Proposition 3.10. Tracing the argument there, the only modification needed is to further compose $B^{2}\mathbb{Z}\rightarrow B^{2}(\mathbb{Z}/2\mathbb{Z})$ with $B^{2}(\mathbb{Z}/2\mathbb{Z})\rightarrow B^{2}R^{\times}$ (the latter map is the twice delooping of $\mathbb{Z}/2\mathbb{Z}=\mathbb{Z}^{\times}\rightarrow R^{\times}$). ∎
By Theorem 1.1, $R$-secondary Maslov data for the canonical Maslov data are given by $R$-spin structures.
Proof of Corollary 1.2.
From Theorem 1.1 we have a canonical Maslov datum for $W$, giving the sheaf of stable categories $\mu sh_{W}$ via Definition 4.25. Also from Theorem 1.1, an $R$-spin structure $\sigma$ on $L$ determines a secondary Maslov datum which, per Lemma 4.21, determines an equivalence $\mu sh_{L,\eta}\simeq\mu sh_{L,\phi_{L}}\simeq loc_{L}$. ∎
Proof of Corollary 1.3.
Without loss of generality, we can assume that $X=L$ is a (conical) complex Lagrangian disk. Endow $W$ with the canonical $R$-Maslov datum. According to Lemma 3.12, there is a canonical choice of secondary grading. So any choice of secondary orientation induces a secondary Maslov datum, and hence, by Corollary 4.28, a microstalk functor $\omega_{p}^{-1}:(\mu sh_{L,o})_{p}\to R-mod$. By Corollary 4.28, $\omega_{p}^{-1}$ acts unambiguously on the set of objects. ∎
We can also generalized Corollary 1.3 to the “constrained” setting (see Section 4.7). Namely, assume that $W$ is an exact complex symplectic manifold and let $L\subset W$ be a (conical, smooth) complex Lagrangian. Fix a coefficient category $\mathcal{C}$ and an anchored collection of subcategories $\{\mathcal{D}_{i}\}$. Endow $W$ with the canonical $(\mathcal{C},\{\mathcal{D}_{i}\})$ grading and assume given a $(\mathcal{C},\{\mathcal{D}_{i}\})$ orientation $o$. According to Lemma 3.12, there is a canonical choice of secondary grading. So any choice of secondary orientation induces a secondary Maslov datum, and hence, by Corollary 4.28, a microstalk functor $\omega_{p}^{-1}:(\mu sh_{L,o})_{p}\to\mathcal{C}$. By Corollary 4.36, the ambiguity of $\omega_{p}^{-1}$ is an element of of $\operatorname{Pic}(\mathcal{C})_{\{\mathcal{D}_{i}\}}$. Hence, given one of the $\mathcal{D}_{i}$, it is meaningful to ask for the image of $\omega_{p}^{-1}$ to be contained in $\mathcal{D}_{i}$. (If $\{\mathcal{D}_{i}\}$ is the total partition of $\mathcal{C}$ into $1$-object subcategories), $\omega_{p}^{-1}:(\mu sh_{L,o})_{p}\to R-mod$ is well-defined as an object, so we recover Corollary 1.3.)
@@ 5.2. Canonical microsheaves and microstalks
Integrability plays no role in the above arguments. That is, for the purpose of defining the canonical microstalk functor $\omega_{p}^{-1}$, it would be enough to assume that $W$ is an exact symplectic manifold endowed with a map $X\to BSp$ lifting the classifying map, and that $L$ is a (conical, smooth) Lagrangian endowed with a complex polarization $L\to LGr_{\mathbb{C}}$.
Our next task is to compute $\omega_{p}^{-1}$ on cotangent bundles of complex manifolds, endowed with the Maslov datum induced by the complex polarization. Concretely, we are interested in the case when $L\subset S^{*}M$ is the conormal to a complex submanifold $N\subset M$ of complex codimension $n$-$\ell$, for $1\leq\ell\leq n$. By Darboux, we may assume without loss of generality that $U=J^{1}\mathbb{R}^{2n}$ and $L$ is the conormal to $\mathbb{R}^{2\ell}\times\{0\}^{2n-2\ell}$. Let $K\subset J^{1}(\mathbb{R}^{2n})$ be the zero section, so that $\psi_{\pi/2}(K)=L$, where $\psi_{\pi/2}(-)$ is defined as in Section 4.6. Henceforth we write $\psi=\psi_{\pi/2}$.
Assume, as above, that $T^{*}M$ is endowed with the canonical $(\mathcal{C},\{\mathcal{D}_{i}\})$ grading and a fixed $(\mathcal{C},\{\mathcal{D}_{i}\})$ orientation $o$. We denote by $\eta$ the corresponding Maslov datum. We let $\phi_{K},\phi_{L}$ denote the canonical polarizations transverse to $K,L$. We denote by $\psi_{*}\eta:=\eta\circ\psi^{-1}$ and $\psi_{*}\phi_{K}:=d\psi(\phi_{K})$ the pushforwards of $\eta$ and $\phi_{K}$ under $\psi$. Observe that $\psi_{*}\phi_{K}:=d\psi(\phi_{K})=\phi_{L}$, by definition of $\psi$.
(18) $\mathcal{C}$$(\mu sh_{K})_{p}$$(\mu sh_{K,\phi_{K}})_{p}$$(\mu sh_{K,\phi_{K}})_{p}=\mathcal{C}$$(\mu sh_{L,\phi_{L}})_{p}$$(\mu sh_{L,\phi_{L}})_{p}=\mathcal{C}$$\mathcal{C}$$(\mu sh_{L})_{p}$$(\mu sh_{L,\phi_{K}})_{p}$$(\mu sh_{L,\phi_{L}})_{p}=\mathcal{C}$$\scriptstyle{\simeq}$$\scriptstyle{[-2\ell]}\scriptstyle{\simeq}$$\scriptstyle{(\psi)_{*}}\scriptstyle{\simeq}\scriptstyle{\simeq}\scriptstyle{\simeq}$$\scriptstyle{\omega_{p}^{-1}}$$\scriptstyle{=}\scriptstyle{\simeq}$$\scriptstyle{=}$$\scriptstyle{[-\ell]}\scriptstyle{\simeq}\scriptstyle$$\simeq$$\scriptstyle{\omega_{p}^{-1}}\scriptstyle{\simeq}$
The dotted arrow means that the rightmost square commutes up to a noncanonical natural transformation (so its effect on objects is unambiguous).
We first explain the meaning of the arrows in (18). The leftmost square is defined as in Lemma 4.13. The map $\mu sh_{K,\phi_{K}}\to\mu sh_{L,\phi_{K}}$ is induced from the contact transformation $\psi_{*}:\mu sh_{T^{*}K,\phi_{K}}\to\mu sh_{T^{*}K,\phi_{K}}$, after stabilizing with $(T^{*}\mathbb{R}^{N},0_{\mathbb{R}^{N}}),N\gg 1$. The identification $\mu sh_{K,\phi_{K}}=\mu sh_{\psi(K),\psi_{*}\phi_{K}}$ and the corresponding top-right commutative square are induced by the contactomorphism $\psi$.
To explain the remaining arrows, recall that $\phi\mapsto(\mu sh_{L,\phi})_{p}$ forms a local system of categories over $LGr=U/O$, which is classified by $U/O\xrightarrow{BJ}BPic(\mathcal{C})$. The postcomposition $U/O\to BPic(\mathcal{C})\to B\mathbb{Z}=U(1)$ is precisely $det^{2}(-)$; hence up to the action of $Pic_{0}(\mathcal{C})$, the monodromy automorphismm of any loop is precisely shifting by the degree of the loop under $det^{2}(-)$.
The canonical microstalk functor $\omega_{p}^{-1}:(\mu sh_{L,\phi_{K}})_{p}\to(\mu sh_{L,\phi_{L}})_{p}=\mathcal{C}$ is realized by choosing a homotopy of polarizations $\phi_{K}\rightsquigarrow\phi_{L}$ through complex Lagrangians, and parallel transporting. Similarly, the arrow $(\mu sh_{L,\phi_{L}})_{p}\to(\mu sh_{L,\phi_{K}})_{p}$ in induced by the homotopy of polarizations $\psi_{t}^{-1}:\psi_{*}\phi_{K}=\phi_{L}\rightsquigarrow\phi_{K}$. Hence the dotted arrow is induced by the loop of polarizations $\phi_{L}\rightsquigarrow\phi_{K}\rightsquigarrow\phi_{L}$, which defines an automorphism $\mathcal{C}=(\mu sh_{L,\phi_{L}})_{p}$. . The first homotopy lies in the kernel of $det^{2}(-)$, while the image of the path $\psi_{t}^{-1}:\psi_{*}\phi_{K}\rightsquigarrow\phi_{K}$ under $det^{2}(-)$ is a loop of degree -$\ell$.
Corollary 5.11.
Let $M,\eta$ be as above. Suppose that $L$ is a smooth Lagrangian disk contained in the the conormal of a complex submanifold $N\subset M$ of complex codimension $\ell$, for $1\leq\ell\leq n$. Let $p\in L$ be a smooth point. Given $A\in\mathcal{C}$, $\omega_{p}^{-1}(A_{N})=A[\ell]$.
Proof.
Without loss of generality, we can assume that $M=\mathbb{R}^{2n}$ and $N=\mathbb{R}^{2\ell}\times\{0\}^{2n-2\ell}$. Let $K=0_{\mathbb{R}^{2n}}$. Then $\omega_{p}^{-1}(A_{K})=A$; by (18), $\omega_{p}^{-1}(A_{N}[-2\ell])=\omega_{p}^{-1}(A_{K})[-\ell]$, which proves the claim. ∎
5.3. Microsheaves on complex contact manifolds and symplectic manifolds